OpenStudy (anonymous):

@umartouseef so here it is . How long would it take for 1.3 C of charge to pass through a point of a wire through which a current of 4.0 mA can pass?

OpenStudy (anonymous):

325 seconds

OpenStudy (anonymous):

how did u do it tho?

OpenStudy (anonymous):

oh and how about this one? A power tool draws 4.5 A of current. If it runs for 1 minute (60 seconds), what is the amount of current that passes through it? do you just multiply? or

OpenStudy (anonymous):

Q=It T=Q/I t=1.3/4*10^-4 t=325seconds

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

well how about this one? A power tool draws 4.5 A of current. If it runs for 1 minute (60 seconds), what is the amount of current that passes through it?

OpenStudy (anonymous):

wil it be 0.075 C?

OpenStudy (anonymous):

the choices are these A 0.075 C B. 4.5 C C. 13.3 C D. 270 C